OpenStudy (anonymous):

How can you find the 1000 letter in " Thursday " ?

OpenStudy (anonymous):

umm...what?

OpenStudy (anonymous):

How do you find the 1000th letter in the word thursday if you keep counting

OpenStudy (zzr0ck3r):

1000/8 what is the remainder?

OpenStudy (anonymous):

125

OpenStudy (anonymous):

it shoulld be less than 8

OpenStudy (anonymous):

no 1000/8 equals 125

OpenStudy (lgbasallote):

the remainder should be less than 8*

OpenStudy (anonymous):

remainder is zero

OpenStudy (zzr0ck3r):

if the remainder was 0 would it not be y?

OpenStudy (anonymous):

it would be y

OpenStudy (zzr0ck3r):

Sorry, I was not trying to give the answer, I just thought you guys were saying its wrong so I wanted to make sure...

OpenStudy (anonymous):

how is it Y

OpenStudy (anonymous):

?

OpenStudy (zzr0ck3r):

imagine we were counting to 16 16/8 has a remainder of 0 notive the 16 letter is y

OpenStudy (zzr0ck3r):

notice*

OpenStudy (anonymous):

Where is the 16 coming from

OpenStudy (zzr0ck3r):

just an example.

OpenStudy (anonymous):

Can you list me the steps pleasee i really need help?

OpenStudy (zzr0ck3r):

i think @KingGeorge is prob doing that.

OpenStudy (kinggeorge):

For example, look at the sequence of letters thursdaythursdaythursdaythursdaythursdaythursday 8 16 24 32 40 48 Notice that every time the letter "y" appears, you see that it is in a position divisible by 8. This is because there are 8 letters in the word "thursday."

OpenStudy (kinggeorge):

It's basically saying the same thing as labeling each number in the set \[\{1,2,3,4,5,6,7,8,9,...,997,998,999,1000\}\]With a letter. You start with the pattern "thursday" and every 8 letters, you repeat that pattern.

OpenStudy (anonymous):

yeah but isnt it 1000/8 which then equals 125 what do you do after that

OpenStudy (kinggeorge):

This means that since your word is 8 letters long, every eighth letter will be "y" starting with the number 8. What you want to do, is find the remainder, and not the quotient. So you have \[{1000}=8\cdot125+0\]So your quotient is 125, and your remainder is 0. Somewhat unintuitively, this means that the 8th letter in "thursday" will be the 1000th letter overall.

OpenStudy (anonymous):

what if i wanna know the 1000th letter of Tuesday

OpenStudy (zzr0ck3r):

1000/7 remainder is what?

OpenStudy (anonymous):

142

OpenStudy (zzr0ck3r):

if its three its the third letter, 4 then 4th, 0 then last

OpenStudy (anonymous):

142.8571429

OpenStudy (kinggeorge):

"tuesday" only has 7 letters. So in this case, you want to find \[1000=7\cdot q+r\]you have \(q=142\). What is \(r\)?

OpenStudy (zzr0ck3r):

we want remainder not dec

OpenStudy (kinggeorge):

\[1000=7\cdot142+r\]\[1000=994+r\]Can you tell me what \(r\) is?

OpenStudy (anonymous):

idk

OpenStudy (kinggeorge):

\[1000-994=994-994+r\]Does this help?

OpenStudy (zzr0ck3r):

1000- 142*7 = r

OpenStudy (zzr0ck3r):

and we will have the rth letter

OpenStudy (anonymous):

Okay thanks